Sean earns $4.50 per hour more than his trainee, Tyrone. During one 8 hour day they earn a total of $284. How much does Sean make per hour?
Thank you

Answer by fractalier(6550)

Let Tyrone's rate of pay be x.
The Sean's rate of pay is x + 4.50.
They both work for eight hours and the total is $284. We write
8x + 8(x + 4.50) = 284
8x + 8x + 36 = 284
16x + 36 = 284
16x = 248
x = 248 / 16
x = 15.50
x + 4.50 = $20.00 per hour